If the regression line between x(score on manual dexterity) and y(units produced per hour)is, y= 19.2+3.0x, then the estimated productivity for a score of 15 will be
Answer by stanbon(57347)

If the regression line between x(score on manual dexterity) and y(units produced per hour)is, y= 19.2+3.0x, then the estimated productivity for a score of 15 will be
----------------------
f(x) = 19.2+3.0x
---
f(15) = 19.2 + 3*15
f(15) = 64.2
================
Cheers,
Stan H.